# Higher-Order Equivariant Neural Networks for Charge Density Prediction in Materials

### v1

@@ Abstract
The calculation of electron density distribution using density functional theory (DFT) in materials and molecules is central to the study of their quantum and macro-scale properties, yet accurate and efficient calculation remains a long-standing challenge in the field of material science. This work introduces ChargE3Net, an E(3)-equivariant graph neural network for predicting electron density in atomic systems. ChargE3Net achieves equivariance through the use of higher-order tensor representations, and directly predicts the charge density at any arbitrary point in the system. We show that our method achieves greater performance than prior work on large and diverse sets of molecules and materials, and scales to larger systems than what is feasible to compute with DFT. Using predicted electron densities as an initialization, we show that fewer self-consistent iterations are required to converge DFT over the default initialization. In addition, we show that non-self-consistent calculations using the predicted electron densities can predict electronic and thermodynamic properties of materials at near-DFT accuracy.
@@ 1 Introduction
Electronic charge density is one of the most fundamental quantities in quantum chemistry and physics, and it is key to accurately modeling molecules and materials at the atomic scale. The Hohenburg-Kohn theorem [1] states that the ground-state charge density contains the information necessary to obtain all ground-state properties of interest. The Kohn-Sham (KS) formulation of density functional theory (DFT) [2] is the most widely used ab initio method for performing electronic structure calculation of molecules or materials. However, KS-DFT is computationally expensive with $O(N^{3})$ complexity, making it infeasible for large-scale quantum calculations on classical computers beyond a few hundred atoms.
More recently, machine learning models have been developed to overcome the computational challenge of ab initio calculations, but they rely on custom-fitted basis functions. Early approaches using machine learning for the prediction of electron density were based on the use of symmetry-adapted Gaussian process regression to predict coefficients for atom-centered basis functions [3, 4]. Coefficients are predicted from a kernel function that expresses the structural similarity and geometric relationship among a target atomic environment and a training set of atomic environments. More recent work has leveraged invariant and equivariant neural networks to predict these coefficients from atomic features. Qiao et al. [5] uses features computed by density-functional tight-binding [6] as inputs for a higher-order equivariant neural network to predict basis set coefficients of small molecules found by self-consistent DFT calculations. Rackers et al. [7] use a higher-order equivariant neural network to predict basis set coefficients directly from atomic graphs. While these approaches are able to achieve high accuracy in specific settings, they are limited by the expressivity of the density fitting basis sets. These atom-centered basis sets are typically defined per-species and are numerically more difficult to converge for solid-state systems compared to plane-wave basis sets, limiting their applicability to molecular systems in vacuum [8].
@@
Alternatively, several methods were proposed to learn electron density directly from a discretized grid of density points. Charge density predictions on grids are agnostic to the basis set used for the ground-truth quantum chemistry calculation and are a natural input format for popular plane-wave DFT codes. By inserting each grid or “probe” point into the atomic graph, charge density prediction can be modeled as a node regression task [9, 10, 11, 12]. However, most of these models only focus on small and specialized atomic systems. Gong et al. [9] use an invariant graph convolutional network [13] to train a charge denisty model on a small dataset of crystalline polymers, demonstrating transferability to unseen structures. Jørgensen and Bhowmik [10] similarly use an invariant graph convolutional network, SchNet [14], demonstrating fast and high accuracy charge density prediction in small molecules, lithium ion battery (LIB) cathode materials and LIB electrolytes. They later demonstrate an improvement in accuracy through the use of an equivariant graph convolutional network, PaiNN [11], through $\mathbb{R}^{3}$ vector representations and rotationally equivariant operations [15]. While these models have been shown to achieve high accuracy when trained and evaluated on small, specialized molecular or material datasets, their expressive power on larger, diverse datasets such as the Materials Project [16] is not yet understood. Furthermore, these models use scalar and/or vector representations to construct rotationally equivariant networks, which could be further improved by incorporating higher-order equivariant features, as has been shown in other tasks, such as the prediction of atomic forces [17].
@@
In this work, we develop a neural network model, ChargE3Net, that uniquely combines graph neural networks, SO(3) equivariant features, and higher-order tensor operations, to accurately predict the electron density of a given atomistic system. Our method enables the use of higher-order equivariant features to better capture the angular resolution and directional information. In a head-to-head comparison, we show that our method achieves greater accuracy than prior work in the prediction of charge density on a diverse sets of density data for both materials and organic molecules. More remarkably, our method outperforms prior models relying on features obtained through quantum mechanical simulation, demonstrating the ability of our approach to learn complex atomic and subatomic interactions from only atomic configurations. Our model trained on over 120K charge density calculations performed by Material Project further demonstrates the expressivity of ChargE3Net, making it applicable to crystalline materials containing any element and of any materials class. To investigate the effect of higher-order representations, we quantify the angular dependence of the charge density at each grid point and show that higher-order representations yield more accurate charge density predictions for materials with high angular variation. We demonstrate the linear time complexity of our model with respect to system size, showing the capability of predicting density on systems with $>10^{4}$ atoms, surpassing what is feasible with ab initio calculations. Lastly, we show that the predicted densities accelerate DFT calculations by reducing the self-consistent steps required to reach a converged solution, and demonstrate the application of using the predicted charge densities to compute a wide range of electronic and thermodynamic properties non-self-consistently, achieving near-DFT accuracies for a majority of materials in a held-out test set.
@@
Figure 1: Overall architecture of the ChargE3Net model. a: Illustration of single charge density probe point, shown in yellow, and its local environment within a periodic atomic system. b: Sub-graph of local environment graph with atom nodes $z_{i}$, $z_{j}$ and probe node $g_{k}$. c: Neural network architecture. Predicted charge density $\hat{\rho}(\vec{g})$ at probe point $\vec{g}_{k}$ is computed from the atomic species $z_{i}$, interatomic displacement vectors $r_{ij}$, and probe-atom displacement vectors $r_{ik}$ through a series of convolution and gate operations. d: $\text{Conv}_{\text{atom}}$ sends higher-order representations bidirectionally to atom nodes. $\text{Conv}_{\text{probe}}$ sends higher-order representations from atoms to probes.
@@ 2.1 Development of higher-order equivariant neural networks for charge density prediction
Given an atomic system (atomic positions and species), we seek to predict the electron charge density $\rho(\vec{r}):\mathbb{R}^{3}\to\mathbb{R}$ at any point within the atomic system. The charge density, along with other properties such as forces, is equivariant under rotation and translation of the atomic system. That is, a rotation or translation of the atomic system in Euclidean space will result in an equivalent rotation or translation to the charge density or forces. Formally this is known as equivariance with respect to E(3), which includes rotations, translations, and reflections in 3D space. This can be achieved in principle by using only invariant scalar features [14], such as inter-atomic distances; however, this prevents the model from using angular information, limiting the accuracy that can be attained [18]. More recent work [15] has introduced equivariance methods using vector $\mathbb{R}^{3}$ features, such as relative atomic positions, to incorporate angular information, and has been shown to improve the performance of charge density prediction [11].
ChargE3Net achieves equivariance through means outlined by tensor field networks [19]. Translation equivariance is achieved by using relative atomic positions; rotation equivariance is achieved by restricting features to be irreducible representations, irreps, of SO(3), which are operated upon by equivariant functions. These features take the form $V^{(\ell,p)}_{cm}$, a dictionary of tensors with keys representing rotation order $\ell\in\{0,1,2,...\}$ and parity $p\in\{-1,1\}$. Each tensor has a channel index $c\in[0,N_{\text{channels}})$ and an index $m\in[-\ell,\ell]$. In this way, the representation at a given $\ell$ and $p$ will have a size of $\mathbb{R}^{N_{\text{channels}}\times(2\ell+1)}$. These representations are combined with the equivariant tensor product operation $\otimes$, using Clebsch–Gordan coefficients $C$ as described in [19] and implemented in e3nn [20]:
@@
where $\ell_{o}$ and $p_{o}$ are given by $|\ell_{1}-\ell_{2}|\leq\ell_{o}\leq|\ell_{1}+\ell_{2}|$ and $p_{o}=p_{1}p_{2}$. We maintain only those representations with $\ell_{o}\leq L$ where $L$ is a maximum allowed rotation order. For ease of notation, we will omit the keys $\ell$, $p$ and indices $c$, $m$ for these tensor representations for the rest of the paper.
@@
As illustrated in Fig. 1(a) and (b), a graph is first constructed with atoms and probe points as vertices, with edges formed via proximity with a cutoff of 4 Å. Edges connecting atoms are unidirectional and edges connecting atoms and probe points are directed pointing from atoms to probe points. Vertex features for atoms $\{A_{i}\}_{i=0}^{N}$ are initialized as a one-hot encoding of the atomic number, represented as a $\ell=0$, $p=1$ tensor with $N_{\rm channels}$ equal to the number of atomic species. Features for probe points $\{P_{k}\}_{k=0}^{M}$ are initialized as a single scalar zero (i.e., $\ell=0$, $p=1$ and $N_{\rm channels}=1$).
ChargE3Net uses a graph neural network (GNN) to update these tensor features, as shown in Fig. 1(c). Within each layer of GNN, the network performs message passing between neighboring atoms and probes, where probe points can only receive messages from neighboring atoms; See Fig. 1(d). The core of the message passing in ChargE3Net is a convolution function that updates the vertex tensor feature with neighboring vertices by the convolution operation defined in Eqn. (4) in Methods. The proposed convolution operation replaces the linear message passing with equivariant operation through tensor product as shown in Eqn. (1) and thus is rotation and translation equivariant; see Methods for more detail. This is then followed by an equivariant Gate activation operation.
ChargE3Net conducts a series of convolution and gate operations. The learned tensor feature associated to each probe point is passed to a post-processing layer where a simple regression decision-head is used to predict the charge density at the corresponding probe location $\hat{\rho}(\vec{g}_{k})$.
@@ 2.2 Performance evaluation
We train and validate ChargE3Net on a diverse set of DFT-computed charge density data, including organic molecules (QM9) [21, 22, 23], nickel manganese cobalt battery cathode materials (NMC) [24], and inorganic materials collected from Materials Project (MP) [25, 16]. The first two datasets have been used in the past for benchmarking charge density models for small molecules and a subset of crystal systems with boundary conditions [11, 5]. To broaden the applicability of the model beyond small, specialized atomic systems, we investigate a more diverse dataset from the MP database, which includes charge densities for over 120K bulk materials. See the Methods section for details of the datasets and the training procedure.
@@
We compare the performance of ChargE3Net to models introduced in OrbNet-Equi [5] and DeepDFT [11]. For QM9 and NMC datasets we use identical training, validation, and test splits as Jørgensen and Bhowmik [11], and report $\epsilon_{\text{mae}}$ computed using the authors publicly released models. We verify these to match the numbers reported in the authors’ original work. For the MP dataset, we train the DeepDFT models using the same experimental settings from the original work. As shown in Table 1, ChargE3Net model significantly outperforms the prior equivariant DeepDFT models [11] on the Materials Project (MP) and QM9 datasets, while achieving similar performance on the NMC dataset. More remarkably, our model achieves a lower $\epsilon_{\text{mae}}$ on the QM9 dataset than OrbNet-Equi [5] model, which leverages additional features based on quantum mechanical calculations while ChargE3Net only uses atomic species and position information. This shows how the model is capable of learning highly informative representations of quantum interactions in a purely data-driven manner.
ChargE3Net’s performance on the MP data demonstrates that it is capable of successfully modelling charge density across the periodic table and a wide variety of material classes that are not present in the other datasets. This suggests ChargE3Net can reproduce a broad range of chemical interactions at the electronic level without incurring the cost of an ab initio calculation. Our results shown in Table 1 also reveal the importance of respecting rotation equivariance in constructing such a general model.
Table 1: Error for each dataset, reported in average $\epsilon_{\text{mae}}$ (%), $\pm$ one standard error. Bold values represent the best $\epsilon_{\text{mae}}$ (%) scores.
@@
2.3 Effects of higher-order representations

### v2

@@ Abstract
The calculation of electron density distribution using density functional theory (DFT) in materials and molecules is central to the study of their quantum and macro-scale properties, yet accurate and efficient calculation remains a long-standing challenge. We introduce ChargE3Net, an E(3)-equivariant graph neural network for predicting electron density in atomic systems. ChargE3Net enables the learning of higher-order equivariant feature to achieve high predictive accuracy and model expressivity. We show that ChargE3Net exceeds the performance of prior work on diverse sets of molecules and materials. When trained on the massive dataset of over 100K materials in the Materials Project database, our model is able to capture the complexity and variability in the data, leading to a significant 26.7% reduction in self-consistent iterations when used to initialize DFT calculations on unseen materials. Furthermore, we show that non-self-consistent DFT calculations using our predicted charge densities yield near-DFT performance on electronic and thermodynamic property prediction at a fraction of the computational cost. Further analysis attributes the greater predictive accuracy to improved modeling of systems with high angular variations. These results illuminate a pathway towards a machine learning-accelerated ab initio calculations for materials discovery.
@@ 1 Introduction
Quantum mechanical interactions of electrons in atoms provide a foundation for understanding and predicting properties of materials [1]. Modern density functional theory (DFT), particularly the Kohn-Sham (KS) formulation of DFT [2], is by far the most widely used method for performing such electronic structure calculations due to its balanced tradeoff between accuracy and computational cost. Because KS-DFT obtains the electronic structure by iteratively solving Kohn-Sham equations and re-calculating the electron charge density until convergence to a self-consistent field (SCF), extensive use of DFT is hindered by the unfavorable $O(N^{3})$ scaling with the system size, making it infeasible for large-scale calculations beyond a few hundred atoms.
While prior works on linear-scaling DFT [3, 4, 5], local coupled cluster methods [6, 7], and orbital-free formulation of DFT [8] have nominally reduced the scaling limit, large-scale calculation is still unobtainable for general problems. In recent years, many efforts have focused on the potential of machine learning (ML) approaches to overcome the quantum scaling limit of DFT. ML models have been developed to map chemical and material systems to properties to enable high-throughput screening [9, 10, 11, 12, 13, 14]. Once trained, ML models are computationally fast, but they are restricted to properties that already exist in a database (typically at the ground state), limiting their applicability to general problems. ML models have also been proven useful for predicting atomic energies and forces for large-scale molecular dynamic simulations [15, 16, 17, 18, 19]. The process generally requires a dedicated computational campaign with DFT to generate high quality data for a subset of materials with a pre-defined potential or temperature range [18, 20, 21, 22, 23, 24, 19]. A more recent work has trained a foundation model across the periodic table for predicting interatomic potentials [25], but it still requires additional DFT data and model finetuning for applications beyond structure relaxation. An alternative approach is to directly predict converged DFT quantities (e.g., electron charge density) to reduce the number of SCF iterations needed to solve KS equations. Since the number of SCF iterations is dependent on the accuracy of the initialized charge density, a fast and accurate charge density prediction may be able to bypass the need for self-consistent iterations, enabling accurate computation of downstream properties at a fraction of the computational cost.
Several works have used ML for charge density prediction. One approach is to expand the charge density with a local atomic basis set and use ML to predict the expansion coefficients [26, 27, 28, 29, 30, 31, 32]. Local atomic orbital basis sets are often parametrized for specific elements. While these models are in principle more efficient, scaling on the order of the number of basis functions, they are restricted to the basis set they are trained on, thus impacting generalizability. Furthermore, any model that attempts to find the parameters of atomic basis sets that best fit grid-based data is fundamentally limited by the expressivity of the basis set [32]. A parallel line of research focuses on the prediction of the Hamiltonian operator [33, 34, 35, 36], however these are also restricted by the choice of basis set, and often not obtainable through DFT codes.
@@
An alternative approach is to learn the electron charge density directly from a discretized grid of density points in 3D Cartesian space. The grid-based approach is agnostic to the basis set (plane-wave or atomic basis) used for the quantum chemistry calculation and are a natural input format for DFT codes [37, 38, 39, 40]. Existing grid-based methods [41, 42, 43, 44, 45] have shown promising results on small molecules and small subsets of material classes. Pioneering works [41, 42] have focused on learning the charge density using graph neural networks invariant to translation, rotation and permutation [9, 46]. Jørgensen and Bhowmik [42] later demonstrate an improvement in accuracy using an equivariant graph neural network, PaiNN [43], through $\mathbb{R}^{3}$ vector representations and rotationally equivariant operations [47]. While these models have been shown to achieve high accuracy on small, specialized molecular or material datasets, no work has demonstrated applicability on larger, diverse datasets across the periodic table. In addition, these models use scalar and/or vector representations to construct rotationally equivariant networks, which could be further improved by incorporating higher-order equivariant features, as has been shown in other tasks, such as the prediction of atomic forces [16]. Finally, limited work has demonstrated the utility of ML-predicted charge densities to initialize DFT calculations, with existing work showing only marginal reduction of SCF steps [48, 44].
@@
In this work, we develop a grid-based model, ChargE3Net, that uniquely combines graph neural networks, SO(3) equivariant features, and higher-order tensor operations, to accurately predict the electron charge density. Our model only uses atomic species and positions as input without the need for additional features. It enables the use of higher-order equivariant features to achieve more accurate performance and higher model expressivity. In a head-to-head comparison with prior grid-based and basis function-based methods, we show that our method not only achieves greater accuracy on existing benchmark data, but also is expressive enough to capture the complexity and variability present in Materials Project (MP) data [49], consisting of more than 100K inorganic materials across the periodic table and for all materials classes. We demonstrate the application of MP-trained ChargE3Net to initialize DFT calculations, resulting in a median of 26.7% reduction in SCF steps on unseen MP data and 28.6% on novel GNoME materials [50]. Both have surpassed the reductions demonstrated by prior work [48, 44]. When evaluating the utility of ChargE3Net in computing a wide range of electronic and thermodynamic properties non-self-consistently, we show a great portion of materials achieving near-DFT accuracy. Further analysis on materials with less-accurate predictions provides insights into a potential cause of the observed inconsistency between self-consistent and non-self-consistent calculations to inform future research. By enabling higher-order equivariant models, we can also investigate the impact of equivariant representations in model performance. We show that higher-order representations yield more accurate charge density predictions for materials with high angular variation. Lastly, We demonstrate the linear time complexity of our model with respect to system size, showing the capability of predicting density on systems with $>10^{4}$ atoms, surpassing what is feasible with ab initio calculations.
@@
Figure 1: Overall architecture of the ChargE3Net model. a Acceleration of DFT property prediction with ChargE3Net charge density initialization. b Illustration of single charge density probe point (center, gray) and its local environment within a periodic atomic system. c Sub-graph of local environment graph with atom nodes $z_{i}$, $z_{j}$ and probe node $g_{k}$. d Neural network architecture. Predicted charge density $\hat{\rho}(\vec{g})$ at probe point $\vec{g}_{k}$ is computed from the atomic species $z_{i}$, interatomic displacement vectors $\vec{r}_{ij}$, and probe-atom displacement vectors $\vec{r}_{ik}$ through a series of convolution and gate operations. e $\text{Conv}_{\text{atom}}$ sends higher-order representations bidirectionally to atom nodes. $\text{Conv}_{\text{probe}}$ sends higher-order representations from atoms to probes.
@@ 2.1 Development of higher-order equivariant neural networks for charge density prediction
The charge density, along with other properties such as forces, is equivariant under rotation and translation of the atomic system. That is, a rotation or translation of the atomic system in Euclidean space will result in an equivalent rotation or translation to the charge density or forces. Formally, this is known as equivariance with respect to E(3), which includes rotations, translations, and reflections in 3D space. This can be achieved in principle by using only invariant scalar features [46], such as inter-atomic distances; however, this prevents the model from using angular information, limiting the accuracy that can be attained [51]. More recent work [47] has introduced equivariance methods using vector $\mathbb{R}^{3}$ features, such as relative atomic positions, to incorporate angular information, and has shown to improve the performance of charge density prediction [43]. We hypothesize that higher-order representations could help improve model performance and accelerate DFT calculations.
ChargE3Net achieves equivariance through means outlined by tensor field networks [52]. Translation equivariance is achieved by using relative atomic positions; rotation equivariance is achieved by restricting features to be irreducible representations, irreps, of SO(3), which are operated upon by equivariant functions. These features take the form $V^{(\ell,p)}_{cm}$, a dictionary of tensors with keys representing rotation order $\ell\in\{0,1,2,...\}$ and parity $p\in\{-1,1\}$. Each tensor has a channel index $c\in[0,N_{\text{channels}})$ and an index $m\in[-\ell,\ell]$. In this way, the representation at a given rotation order $\ell$ and parity $p$ would have a size of $\mathbb{R}^{N_{\text{channels}}\times(2\ell+1)}$. These representations are combined with the equivariant tensor product operation $\otimes$, using Clebsch–Gordan coefficients $C$ [52] as implemented in e3nn [53]:
@@
where $\ell_{o}$ and $p_{o}$ are given by $|\ell_{1}-\ell_{2}|\leq\ell_{o}\leq|\ell_{1}+\ell_{2}|$ and $p_{o}=p_{1}p_{2}$. We maintain only those representations with $\ell_{o}\leq L$ where $L$ is a maximum allowed rotation order. For ease of notation, we will omit the keys $\ell$, $p$ and indices $c$, $m$ for these tensor representations for the rest of the paper. Unless otherwise indicated, we use a $L=4$ maximum rotation order for ChargE3Net.
@@
As illustrated in Fig. 1b and 1c, a graph is first constructed with atoms and probe points as vertices, with edges formed via proximity with a cutoff of 4 Å. Edges connecting atoms are unidirectional and edges connecting atoms and probe points are directed pointing from atoms to probe points. Vertex features for atoms $\{A_{i}\}_{i=0}^{N}$ are initialized as a one-hot encoding of the atomic number, represented as a $\ell=0$, $p=1$ tensor with $N_{\rm channels}$ equal to the number of atomic species. Features for probe points $\{P_{k}\}_{k=0}^{M}$ are initialized as a single scalar zero (i.e., $\ell=0$, $p=1$ and $N_{\rm channels}=1$).
ChargE3Net uses a graph neural network (GNN) to update these tensor features (Fig. 1d). Within each layer of GNN, the network performs message passing between neighboring atoms and probes, where probe points can only receive messages from neighboring atoms (Fig. 1e). The core of the message passing in ChargE3Net is a convolution function that updates the vertex tensor feature with neighboring vertices by the convolution operation (Eqn. (4) in Methods). Since the proposed convolution operation replaces the linear message passing with equivariant operation through the tensor product defined in Eqn. (1), the operation is rotation and translation equivariant; see Methods for more detail.
ChargE3Net conducts a series of equivariant convolution and gate operations to obtain representations for each atom, with an addition series of equivariant convolution and gate operations to obtain representations for each probe point. The learned tensor feature associated to each probe point is passed to a post-processing layer where a simple regression is used to predict the charge density at the corresponding probe location $\hat{\rho}(\vec{g}_{k})$.
@@ 2.2 Performance evaluation
We train and validate ChargE3Net on a diverse set of DFT-computed charge density data, including organic molecules (QM9) [54, 55, 56], nickel manganese cobalt battery cathode materials (NMC) [57], and inorganic materials collected from Materials Project (MP) [58, 49]. The first two datasets have been used in the past for benchmarking charge density models for small molecules and a subset of crystal systems with boundary conditions [43, 30]. To develop a charge density model for the entire periodic table, we investigate one of the largest open databases of DFT, the Materials Project (MP) database. It includes charge densities for over 100K bulk materials with diverse compositions and structures (Supplementary Fig. A.1-A.3 and Methods).
@@
Table 1: Model performance for charge density prediction
@@
All values are presented as an average $\epsilon_{\text{mae}}$ (%) and $\pm$ one standard error. Bold values represent the best $\epsilon_{\text{mae}}$ (%) scores. $\epsilon_{\text{mae}}$ reported for MP include both magnetic and non-magnetic materials.
Table 1 compares the performance of ChargE3Net to models introduced in OrbNet-Equi [30] and DeepDFT [43]. For QM9 and NMC datasets, we use identical training, validation, and test splits as Jørgensen and Bhowmik [43]. All reported $\epsilon_{\text{mae}}$ are computed using publicly released models and verified to match the numbers reported in the original work. For the MP dataset, we train the DeepDFT models using the same experimental settings from the original work (see Supplementary section for more detail). We observe that all equivariant models perform better than invariant models, demonstrating the importance of respecting rotation equivariance. Moreover, ChargE3Net model significantly outperforms prior equivariant DeepDFT models [43] on the Materials Project (MP) and QM9 datasets, while achieving similar performance on the NMC dataset. This suggests that the inclusion of higher-order representations can improve model performance. More remarkably, our model achieves a lower $\epsilon_{\text{mae}}$ on the QM9 dataset than OrbNet-Equi [30] model, which leverages additional features based on quantum mechanical calculations while ChargE3Net only uses atomic species and position information. This shows how the model is capable of learning highly informative representations of quantum interactions in a purely data-driven approach.
ChargE3Net’s performance on the MP data demonstrates that it is capable of successfully modelling charge density across the periodic table and a wide variety of material classes that are not present in other datasets. When we compare the prediction performance with respect to the number of unique species in a material, the MP-trained ChargE3Net is capable of predicting charge density in systems with many unique species, without a significant change in error distribution as number of species increases (Supplementary Fig. A.4). This suggests ChargE3Net can reproduce a broad range of chemical interactions at the electronic level without incurring the cost of an ab initio calculation.
2.3 Using predicted charge densities to accelerate DFT
DFT calculations require an initial charge density and wavefunction prior to iteratively solving the Kohn-Sham equations until self-consistency [59]. Charge density initializations that closely match the ground-state charge density often save compute time by reducing the number of self-consistent field (SCF) steps needed to find a self-consistent solution. Many schemes for initializing the charge density exist in the literature, but the most common one is the superposition of atomic densities (SAD), where the initial charge density is computed by summing over the atomic charge density of each atom in the system [60].
We demonstrate that charge densities predicted by ChargE3Net reduce the number of SCF steps, thereby accelerating DFT calculations, on unseen non-magnetic materials (Fig. 2a, b) compared to a SAD initialization. The median number of SCF steps needed to converge a single-point calculation self-consistently for a non-magnetic material starting from the SAD initialization is 15, while the same starting from a ChargE3Net initialization is 11, representing a 26.7% decrease in SCF steps. As an upper-bound baseline, we also include the distribution of the number of SCF steps needed to converge a calculation starting from a converged (i.e. ground-truth) charge density (SC), showing that our model still has room for improvement. About 96% of calculations done with a predicted charge density ran with fewer SCF steps than initializing with SAD, and the compute time improvement is robust to a wide range of $\epsilon_{mae}$ (Supplementary Fig. A.5).
Figure 2: DFT calculations with ChargE3Net. Comparison of DFT calculations initialized from superposition of atomic densities (SAD), ChargE3Net predictions, and ground-truth densities from existing self-consistent calculations (SC). a Average number of SCF steps required to converge the total energy to $5\times 10^{-5}$ eV/atom for non-magnetic materials in the MP test set, with respect to the number of unique species within each material. Error bar shows $\pm$ one standard error. b Number of steps required for self-consistent (SC) and non-self-consistent (NSC) calculations. Box-plots show median as center line, upper and lower quartiles as box limits, 1.5x interquartile range as whiskers, and outliers as points. c Non-self-consistent physical property calculations for total energy, atomic forces, band energies, and band gaps. Dashed lines indicate chemical accuracy (1 meV/atom for per-atom energies, 0.03 eV/Å  for forces, and 43 meV for band energies and band gap). Errors of exactly zero are replaced with $10^{-8}$ for visualization purposes. Interior of violin-plots shows median as center line, upper and lower quartiles as box limits, and 1.5x interquartile range as whiskers. d Percentage of materials achieving chemical accuracy for non-self-consistent calculations.
To evaluate how well ChargE3Net’s predicted charge densities reproduce physical properties, we compute the energy, forces, band energies, and bandgaps of the test set non-self-consistently (Fig. 2c, d). 40% of materials in the test set have per-atom energy errors less than 1 meV/atom. 70% of materials have forces less than 0.03 eV/Å, a common threshold for structural relaxations; 15% of the test set have force errors of 0.0 (Supplementary Fig. A.6). All of the structures with perfect agreement have self-consistent forces of 0.0 eV/Å for each atom and Cartesian component, possibly because of cancellation of charge density errors due to symmetry. 45% and 76% of materials recover band energies at each k-point and the direct band gap within chemical accuracy (43 meV), respectively. 45% of the band gaps obtained non-self-consistently exactly matched those of self-consistent calculations (Supplementary Fig. A.6). Similar to the force data described above, calculations with perfect agreement all have a self-consistent band gap of 0 eV, indicating the model is able to readily identify metallic materials. We note that our use of chemical accuracy for band gaps is a more stringent criterion than experimental measurement errors for well-studied semiconductors [61]. In addition, non-self-consistent calculations using the Materials Project parameters as done with this test set have an additional source of error (Supplementary Fig. A.7 and Methods) as a result of an incorrect calculation parameter set by the Materials Project. This error could explain why not all NSC calculations computed on a self-consistent charge density achieve chemical accuracy (see Fig. 2d and Fig. A.8).
To examine generalization capabilities of ChargE3Net on out-of-distribution data, we analyze a random subset of 1924 materials published by GNoME [50] with 5 or more unique species to differentiate it from the MP distribution. The average $\epsilon_{mae}$ found for this set is 0.632%, which is higher than that of the MP test set for ChargE3Net but lower than either DeepDFT variant trained on the Materials Project (Table 1). We conduct a similar experiment to the MP test set above with non-self-consistent DFT calculations on 1510 non-magnetic materials. The median energy error is 4.3 meV/atom for the GNoME set, whereas the same is 1.8 meV/atom for the MP test set. Only 8.4% of materials in GNoME computed non-self-consistently from the ChargE3Net charge density obtained the total energy within chemical accuracy (<1 meV/atom), though nearly all self-consistent calculations meet this criteria (Supplementary Fig. A.9). The median number of SCF steps needed for self-consistency falls from 14 to 10 steps when initializing with ChargE3Net, a 28.6% drop (Supplementary Fig. A.10) and a similar acceleration found for the Materials Project test set. Moreover, 91% of calculations saw a decreased number of SCF steps to reach convergence. While further work is needed to improve the generalizability of ChargE3Net for non-self-consistent property predictions, self-consistent calculations on out-of-distribution materials are still accelerated when initialized by the predicted charge density.
In addition to the model’s robustness to materials outside its training set, we show that ChargE3Net can predict the charge densities of materials after small atomic perturbations with minor increase to the charge density error (Supplementary Fig. A.11). This allows us to accurately obtain phonon properties, which may rely on finite differences as the main method to obtain approximations to the local potential energy surface, from non-self-consistent calculations. We compute the constant-volume heat capacity ($C_{v}$) under the harmonic approximation for two zeolites, which are not included in the Materials Project dataset, and show that $C_{v}(T)$ obtained closely match those obtained by self-consistent calculations (Supplementary Section A.2). As another example of transferability, we compute the electronic bandstructure of the Si(111)-2x1 surface non-self-consistently from the ChargE3Net prediction and show that it almost exactly matches that of a self-consistent calculation (Supplementary Section A.3).
2.4 Higher-order representations lead to improved model accuracy